The sum of the ages of Ruth and her mother is 75 years. The difference in thier ages is 29 years. How old is each?
r+m=75?
m-r=29?

Answer by stanbon(75887)

The sum of the ages of Ruth and her mother is 75 years. The difference in thier ages is 29 years. How old is each?
r+m=75?
m-r=29?
-----------
Add the two equations to get:
2m = 104
m = 52 (mother's age)
---
Substitute to solve for "r":
r + 52 = 75
r = 23 (Ruth's age)
=====================
Cheers,
Stan H.